OpenStudy (anonymous):

Perform the indicated operation. 2/3 + 3/8

OpenStudy (abdullahm):

make sure that both fractions have the same denominator.

OpenStudy (abdullahm):

What is the least common multiple of 3 and 8?

OpenStudy (anonymous):

2?

OpenStudy (abdullahm):

Do you know what multiples are?

OpenStudy (anonymous):

no..

OpenStudy (abdullahm):

Multiples of 3: 3, 6, 9, 12, 15, 18, 21... and so on

OpenStudy (abdullahm):

What are the multiples of 8?

OpenStudy (anonymous):

16 24 and 32?

OpenStudy (abdullahm):

so the least common multiple would be 24

OpenStudy (anonymous):

why wouldnt it be 16? because thats the lowest other than 8

OpenStudy (abdullahm):

But is 16 divisible by 3? No.

OpenStudy (anonymous):

oh okay i get that

OpenStudy (abdullahm):

\(\sf\Large\left(\frac{2}{3} \times\frac{8}{8}\right)+\left(\frac{3}{8}\times\frac{3}{3}\right)=?\)

OpenStudy (abdullahm):

Hint: \(\sf\Large\frac{a}{b}\times\frac{c}{d}=\frac{a~\times~ b}{c ~\times~ d}\)

OpenStudy (anonymous):

16/24 + 9/24

OpenStudy (anonymous):

@AbdullahM so 25/24

OpenStudy (abdullahm):

correct and that is 1 1/24

OpenStudy (anonymous):

@AbdullahM how do you get it like that though?

OpenStudy (abdullahm):

because \(\sf\Large\frac{25}{24} =\frac{24+1}{24}=\frac{24}{24}+\frac{1}{24}=1\frac{1}{24}\)

OpenStudy (anonymous):

@AbdullahM oh okay i get that, thanks a lot